# Mixed Hodge Modules and Canonical Perverse Extensions for Multi-Node Conifold Degenerations

### 07 Apr 2026 v1

@@ 9.3. Further directions
The theorem package established here resolves the finite multi-node ordinary double point case, but it also points directly to a number of natural next problems.
(1) Stronger comparison of extension classes.
@@
The present paper now relates the mixed-Hodge-module extension, the realized perverse extension, and the induced extension on hypercohomology. A natural next step is to strengthen this further into a more explicit comparison theorem among extension classes in $MHM(X_{0})$, in $\mathrm{Perv}(X_{0};\mathbb{Q})$, and in the category of mixed Hodge structures, with sharper control of the maps between these extension spaces and of the corresponding nodewise generators.
(2) Hodge-theoretic lifts of the nodewise Ext generators.
@@
Section 5 isolates distinguished nodewise extension classes on the perverse side. A natural next problem is to construct and understand corresponding classes directly in
$\mathrm{Ext}^{1}_{MHM(X_{0})}\!\left(i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),IC^{H}_{X_{0}}\right)$
and to compare them functorially with the perverse and hypercohomological extension classes. Such a refinement would sharpen the rigidity theory and make the Hodge-theoretic content of the global extension more explicit.
(3) Arbitrary stratified singular loci.
@@
The most immediate geometric extension is to pass beyond finite collections of ordinary double points and treat degenerations whose singular locus has higher-dimensional strata. At the perverse level, the corrected object in such a setting is expected to differ from the intersection complex by a constructible singular contribution supported on the singular locus. The mixed-Hodge-module analogue would require replacing the finite direct sum of point-supported rank-one objects by a more general singular quotient, together with a corresponding extension of the divisor-gluing construction carried out here.
(4) Kähler-package questions for $\mathbb{H}^{*}(X_{0},\mathcal{P}^{H})$.
@@
Once the mixed-Hodge-module refinement $\mathcal{P}^{H}$ is available, it becomes meaningful to ask whether its hypercohomology satisfies analogues of the classical Kähler-package properties known for intersection cohomology: duality, Lefschetz-type isomorphisms, and Hodge–Riemann bilinear relations. The present paper establishes the Hodge-theoretic object to which such questions may now be addressed, but does not attempt to resolve them.
(5) Quiver, schober, and wall-crossing refinements.
@@
The finite-node extension data already suggests a quiver-theoretic organization of the localized node contributions and their coupling to the bulk sector. Likewise, the relation of the corrected perverse object to spherical-monodromy phenomena suggests that a categorified schober-theoretic version of the present construction should exist. The mixed-Hodge-module refinement proved here is the natural precursor of those higher-categorical structures.
(6) Full LMHS and coefficient interpretation.
@@
The present paper identifies the point-supported quotient as the source of the finite vanishing sector in the limiting mixed Hodge structure. A natural next question is to understand more fully the coefficients with which the global corrected class is assembled from the nodewise sectors and to give those coefficients a direct Hodge-theoretic and limiting-mixed-Hodge-theoretic interpretation.
Taken together, these directions show that the finite multi-node ordinary double point theorem proved here is not the end of the story, but the first fully internal Hodge-theoretic foundation for a broader program linking nearby cycles, mixed Hodge modules, canonical perverse extensions, extension spaces, quiver-type structures, and physical wall-crossing phenomena.

### 09 Apr 2026 v2

@@ 9.3. Further directions
The finite multi-node ordinary double point case treated here suggests several natural next problems.
  1. (1)
@@
    Sharper comparison of extension classes. The present paper relates the mixed-Hodge-module extension, its realized perverse extension, and the induced extension on hypercohomology. A natural next step is to strengthen this into a more explicit comparison theorem among extension classes in $MHM(X_{0})$, in $Perv(X_{0};\mathbb{Q})$, and in the category of mixed Hodge structures.
  2. (2)
@@
    Hodge-theoretic lifts of nodewise generators. Section 5 isolates distinguished nodewise extension classes on the perverse side. A natural problem is to construct corresponding classes directly in
    $\mathrm{Ext}^{1}_{MHM(X_{0})}\bigl(i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),IC^{H}_{X_{0}}\bigr)$
    and compare them functorially with the perverse and hypercohomological classes.
  3. (3)
@@
    Beyond finite ordinary double points. The next geometric extension is to pass from a finite set of ordinary double points to more general stratified singular loci. In that setting the point-supported singular quotient used here would have to be replaced by a more general singular contribution together with a corresponding extension of the divisor-gluing construction.
  4. (4)
@@
    Kähler-package questions for $H^{\ast}(X_{0},\mathcal{P}^{H})$. Once the mixed-Hodge-module refinement $\mathcal{P}^{H}$ is available, it becomes meaningful to ask whether its hypercohomology satisfies analogues of the classical package for intersection cohomology, including Lefschetz-type statements and Hodge–Riemann-type relations.
  5. (5)
@@
    Quiver, schober, and wall-crossing refinements. The finite-node extension data already suggests a quiver-type organization of the localized node contributions and their coupling to the bulk sector. Likewise, the relation of the corrected perverse object to spherical-monodromy phenomena suggests a categorified refinement. The mixed-Hodge-module construction proved here is the natural precursor to those higher-categorical structures.
  6. (6)
@@
    Full LMHS interpretation of the nodewise coefficients. The present paper identifies the point-supported quotient as the source of the finite vanishing sector in the limiting mixed Hodge structure. A natural next question is to understand more explicitly the coefficients with which the global corrected class is assembled from the nodewise sectors and to give those coefficients a direct Hodge-theoretic interpretation.
These directions all build on the same point: the corrected extension is now available as an object internal to $MHM(X_{0})$, together with its realization and its induced hypercohomological comparison. That is the foundational result established here.